OpenStudy (anonymous):

Will give a medal for the correct answer! Find three consecutive odd integers such that the sum of the largest and twice the smallest is 25.

ganeshie8 (ganeshie8):

i wud start by testing few :- 1, 3, 5 5+2x1 = 7 too less, discard

OpenStudy (anonymous):

2x + x + 2 = 25 2x + x + 4 = 25 2x + 2x + 4 = 25 2x + 2x + 2 = 25 are the choices <_<

ganeshie8 (ganeshie8):

oh they want u setup an equation to solve it. they dont want u solve it.